Write an equation of a line that is perpindicular to y=3x-3 and runs through (10,-3).
Answer by mananth(16946)

1 y = 3 x -3
Divide by 1
y = 3 x + -3
Compare this equation with y=mx+b
slope m = 3

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1


m= - 1/3 ,point ( 10 , -3 )
Find b by plugging the values of m & the point in
y=mx+b
-3 = -3.33 + b
b= 0.33
m= - 1/3
The required equation isy = - 1/ 3 x + 1/ 3
m.ananth@hotmail.ca